Dian S. answered 02/01/21
AP Statistics Tutor with Graduate Level teaching experience
Each trial has two outcomes: gets a hit, or does not get a hit. Assuming each hit is independent, p stays constant at 0.177. n is given to be 6
X is distributed as a Binomial with n=6, p-0.177
P(X>=4) = 1-P(X<=3) = 1-binomcdf(6,0.177,3) = 0.0109